# DecompDreamer: Advancing Structured 3D Asset Generation with Multi-Object Decomposition and Gaussian Splatting

### v1

@@ 4 Method
Our objective is to generate high-quality 3D assets for complex text prompts that describe intricate inter-object relationships. Figure 3 shows an overview, with details described below. In section  4.1, we discuss rectified flow based distillation. In section  4.2, we discuss the structured decomposition of compositional text into objects and their relationship via VLM. Finally, in section  4.3, we discuss the progressive optimization process, transitioning from relationship modeling to object refinement for high-fidelity compositional 3D generation.
4.1 Flow Matching Distillation (FMD)
Rectified flow models employ a linear interpolation forward process given by $z_{t}=(1-\sigma_{t})z+\sigma_{t}\epsilon$, or equivalently, $z_{t}=z-\sigma_{t}(\epsilon-z)$. This process interpolates between the data $z$ and the noise $\epsilon$ using a time-dependent weighting function $\sigma_{t}$, which is sampled from a uniform distribution $\mathcal{U}(0,1)$. For simplicity, we denote $(\epsilon-z)$ by $u(z,\epsilon)$, representing the actual vector field.
FMD optimizes a 3D representation $\theta$ using a pre-trained rectified flow-based model $\phi$. The optimization, which closely aligns with the SDS gradient, is formulated as
$\nabla_{\theta}\mathcal{L}_{\text{FMD}}(z;y)=\mathbb{E}_{t,\epsilon}\Big{[}\sigma_{t}\big{(}v_{\phi}(z_{t},t,y)-u(z,\epsilon)\big{)}\frac{\partial z}{\partial\theta}\Big{]},$ (2)
where $v_{\phi}(.)$ is the vector field predicted by the model $\phi$, and where $z_{t}$, $t$, and $y$ correspond to the noisy image, the timestep, and the prompt embedding, respectively. This optimization minimizes the discrepancy between the predicted and the actual vector fields, thereby guiding the 3D representation $\theta$ to produce renderings that are consistent with the given text prompt.
@@ 4.2 Structured Decomposition with VLMs
Before designing a complex 3D asset, an artist typically begins with careful planning and wireframing the overall structure. This involves determining the placement, size, and orientation of individual objects within the scene. Inspired by this process, we leverage a VLM, specifically ChatGPT-4o [30], to structure our 3D representation. Given a text prompt $t^{g}$, the VLM generates an image $I$ along with a corresponding scene graph $G=(V,E)$, where $V$ represents the set of objects and $E$ captures the inter-object relationships. We explicitly instruct the VLM to generate a connected graph, ensuring each node has at least one edge.
The objects identified by the VLM from the text prompt $t^{g}$ are denoted as $\{O_{i}\}_{i=1}^{n}$, where $n$ is the total number of objects extracted. These objects serve as the nodes $V=\{v_{i}\}_{i=1}^{n}$ in the scene graph $G$. The VLM extracts additional descriptions in the form of attributes $\{a_{ij}\}_{j=1}^{m}$ from the text prompt $t^{g}$ and associates them with their corresponding nodes in the scene graph. By concatenating an object $O_{i}$ with its attributes, we construct an object prompt $t^{o}_{i}$. For example, in Figure 3, the combination of the object and its attributes yields the object prompt: “A milkshake with two straws”. Similarly, the relationships between object pairs $(O_{i},O_{j})$, as described in $t^{g}$, are mapped to edges $e_{ij}$ within the scene graph $G$. These relationships capture the contextual interaction between the objects. To represent these interactions, we construct edge prompts $t^{e}_{ij}$ by concatenating the involved objects and the relationship attribute. By integrating both object-level and relational details through object and edge prompts, the scene graph serves as a comprehensive blueprint, effectively guiding the 3D representation process.
Additionally, we pass the generated image $I$ and scene graph $G$ to the VLM to infer spatial and geometric properties such as center coordinates, size, and orientation of each object. These properties collectively define the spatial arrangement of the 3D representation.
Coarse Initialization. Our primary objective is to accurately capture complex object relationships while ensuring proper disentanglement of individual elements within the scene. To achieve this, we initialize the scene using 3D Gaussians, where each Gaussian is assigned to a specific object $O_{i}$. The Gaussians are initialized from point clouds generated by Point-E [28] based on object-specific prompts. Each object’s point cloud is aligned using its inferred center, scale, and orientation to maintain scene consistency. The full scene representation is obtained by aggregating all aligned point clouds, forming a structured initialization for further optimization. For efficient disentangled optimization, we maintain a mapping that tracks Gaussians corresponding to each object, enabling independent refinement of properties such as mean, covariance, and opacity. Further details, including transformation equations and formal definitions, can be found in Section A.3 of the Appendix.
4.3 Progressive FMD Optimization
@@ 4.3.1 Ensuring Coherent Inter-Object Interactions
Joint Optimization of Related Objects: Given an edge prompt $t^{e}_{ij}$ describing the relationship between objects $O_{i}$ and $O_{j}$, we jointly optimize the Gaussians corresponding to both objects. This ensures that the spatial and semantic relationships between the objects are preserved in the 3D scene. The FMD loss for this joint optimization is :
$\mathcal{L}_{joint}(O_{i},O_{j},y^{e}_{ij})=\nabla_{\mathcal{G}_{o_{i}},\mathcal{G}_{o_{j}}}\mathcal{L}_{\text{FMD}}(z;y^{e}_{ij}),$ (3)
where, $\mathcal{L}_{\text{FMD}}$ is defined in (2), $\mathcal{G}_{o_{i}}$ and $\mathcal{G}_{o_{j}}$ represent the Gaussians for objects $O_{i}$ and $O_{j}$, respectively, and $y^{e}_{ij}$ is the text embedding corresponding to the edge prompt $t^{e}_{ij}$.
Targeted Optimization for a Single Object: While joint optimization ensures cohesive inter-object relationships, it can sometimes result in undesirable blending, where parts of one object flow into another. To mitigate this, we employ targeted optimization based on the edge prompt, focusing the optimization solely on one object while keeping the other fixed. The goal is to adjust the selected object such that it adheres to the overall relationship without affecting the structural integrity of the other object. The FMD loss for targeted optimization is formulated as:
$\mathcal{L}_{target}(O_{i},y^{e}_{ij})=\nabla_{\mathcal{G}_{o_{i}}}\mathcal{L}_{\text{FMD}}(z;y^{e}_{ij}),$ (4)
@@
Scene-Aware Gaussian Translation: When modeling inter-object relationships, it is important to account for errors in object center predictions by the VLM. To address this, we introduce object-specific translation parameters $T^{o}_{i}=(x_{i},y_{i},z_{i})$ for each object $O_{i}$. Before optimizing edges using (3) or (4), Gaussians are translated using these parameters to correct misalignments and ensure accurate spatial relationships. The transformation is applied as:
$\mathcal{G}_{o_{i}}\leftarrow\mathcal{G}_{o_{i}}+T^{o}_{i},\quad\mathcal{G}_{o_{j}}\leftarrow\mathcal{G}_{o_{j}}+T^{o}_{j}$
where $T^{o}_{i}=(x_{i},y_{i},z_{i})$ and $T^{o}_{j}=(x_{j},y_{j},z_{j})$ represent the translation adjustments for objects $O_{i}$ and $O_{j}$, respectively.
@@ 4.3.2 View-Aware Disentangled Object Optimization
DecompDreamer promotes object disentanglement by assigning each Gaussian to a single object, ensuring clear boundaries. However, solely optimizing Gaussians based on inter-object relationships can lead to object blending and suboptimal refinement of individual objects. This section discusses optimizing objects based on attributes and enhancing disentanglement using negative prompts.
Optimizing Objects within Compositional Environment: A naïve approach would apply SDS-based optimization directly using object prompts, but in a multi-object scene, relative object orientation significantly impacts perception. DreamFusion [32] demonstrated that view-dependent text prompts improve SDS-based optimization by aligning prompts with the camera’s viewpoint. Let $t$ be the base text prompt. The view-dependent embedding $y_{a}$ is generated using $E(t,a)$, which integrates $t$ with the azimuth angle $a$ to produce a view-aware embedding. This method interpolates between front, side, and back views based on $a$, aligning text prompts with the camera’s viewpoint. This alignment enhances orientation-specific details, leading to more coherent 3D outputs and has become standard in optimization-based 3D generation. However, global view-dependent prompts fail in multi-object scenes where a single camera view may not accurately represent all objects. To address this, we introduce object-view-dependent text prompts, ensuring each object’s orientation aligns with its correct perspective. First, each object $O_{i}$ is rendered at a canonical front-facing orientation ($0^{\circ}$ azimuth) to establish a reference view. The VLM then estimates the angular difference $\phi_{i}$ between the current view and the object’s front view, correcting the azimuth as $a_{i}=a-\phi_{i}$ (a sample VLM workflow is provided in Section A.5 of the appendix). The adjusted azimuth $a_{i}$ is incorporated into the object’s text prompt, leading to the object-specific FMD loss:
$\mathcal{L}_{obj}(O_{i},y^{o}_{a_{i}})=\nabla_{\mathcal{G}_{o_{i}}}\mathcal{L}_{\text{FMD}}(z;y^{o}_{a_{i}}),$ (5)
where $y^{o}_{a_{i}}=E(t^{o}_{i},a)$. This adjustment ensures more coherent 3D reconstructions by aligning object prompts with their true viewpoints.
Negative Prompt for Disentangled Object: As discussed earlier, optimizing inter-object relationships can lead to object blending, where one object undesirably flows into another. To mitigate this, we incorporate negative prompts during individual object optimization. The negative prompt for an object $O_{i}$ is constructed by aggregating all objects directly connected to $O_{i}$ in the scene graph, ensuring that the model explicitly avoids incorporating features from these related objects. This approach helps maintain clear object boundaries and preserves structural integrity within the scene.
Figure 4: Qualitative comparison between the proposed DecompDreamer and state-of-the-art generators, including GALA3D [43], GraphDreamer [13], MVDream [36], LucidDreamer [20] and Magic3D [21]. More results can be found in Section A.6 of the appendix.
@@ 4.3.3 Overall Optimization
After initialization, we begin by modeling inter-object relationships before gradually shifting focus toward individual objects. The optimization process is divided into two stages:
1. Stage One: We emphasize joint modeling of inter-object relationships. For each edge in the scene graph, we apply joint optimization (3), while progressively increasing the weight of individual object optimization (5).
2. Stage Two: The focus shifts toward refining individual objects while maintaining inter-object consistency. Here, we perform targeted optimization for each object (4) along with individual object optimization (5). This transition allows fine-grained details to be added while preserving relationships established earlier. The effective FMD loss is:
$\mathcal{L}_{\text{loss}}=\begin{cases}\begin{array}[]{l}\mathcal{L}_{joint}(O_{i},O_{j},y^{e}_{ij})\\ +\lambda.\left(\frac{t}{T}\right)^{2}\sum\limits_{O\in\{O_{i},O_{j}\}}\mathcal{L}_{obj}(O,y^{o}_{a_{o}}),\\ \end{array}&t<\gamma.T\\[8.0pt] \\ \begin{array}[]{l}\sum\limits_{O\in\{O_{i},O_{j}\}}(\mathcal{L}_{target}(O,y^{e}_{ij})\\ +\mathcal{L}_{obj}(O,y^{o}_{a_{o}})),\end{array}&t>\gamma.T\end{cases}$ (6)
where $(i,j)\in E$, $t$ represents the current iteration, $T$ is the total number of iterations, and $\gamma$ controls the transition point between joint and targeted optimization. After a complete cycle of optimizing each edge, we refine the entire scene based on the original text prompt. The full pipeline, including algorithmic details, is provided in Section A.1 of the Appendix.
@@ 5 Experiments
Figure 5: (a) Qualitative comparison of object disentanglement in GALA3D, GraphDreamer, and DecompDreamer. Each method displays the full 3D object (left) and its individual components (right). (b) Ablation study on with and without the azimuth adjustment for object-view aware object optimization.
Figure 6: Results of an ablation study on DecompDreamer with the text prompt: “A knight in full shining armor galloping on a brown horse, wearing a blue hat, and holding a sword”. Additional ablation studies can be found in Figure  8 of the appendix.
Figure 7: Qualitative comparison of GALA3D and GraphDreamer against DecompDreamer with SD2.1 and SD3.5 as backbones.
We implement our guidance model using the publicly available diffusion model, stabilityai/stable-diffusion-3.5-medium, with a guidance scale of 7.5. For relationship optimization, we adopt the view-dependent prompting technique from DreamFusion, while for object optimization, we introduce object-view-dependent text prompts, where the corrected azimuth angle is estimated by the VLM. Our 3D Gaussian Splatting rendering pipeline is implemented in PyTorch, with the total number of iterations determined by the number of objects, allocating 1500 iterations per object. Each object’s Gaussian representation is initialized with 80K points. Generating a 3D compositional asset with two objects typically requires 90 minutes on an A100 GPU, averaging 45 minutes per object. Further implementation details and hyperparameter settings can be found in Section A.2 of the Appendix . We qualitatively and quantitatively compare DecompDreamer with state-of-the-art text-to-3D generators on compositional text prompts in Section 5.1. We present our ablation studies in Section 5.2 on significance of various component of DecompDreamer.
@@ 5.1 Comparison to State-of-the-art Methods
We qualitatively and quantitatively evaluate DecompDreamer against state-of-the-art (SOTA) text-to-3D models [43, 13, 36, 20, 21] using complex compositional prompts. Another notable work, CompGS [14], also achieves SOTA results, however, we are unable to compare with them as their code is not available at the time of writing this paper. As shown in Figures 1, 4 and 15, DecompDreamer effectively models intricate inter-object relationships while scaling to a large number of objects. GALA3D is capable of generating large-scale environments (e.g., a bedroom) but struggles with modeling relationships, even in simpler cases like a knight riding a horse. This limitation arises because it does not explicitly model individual relationships, relying instead on global scene optimization. GraphDreamer jointly optimizes objects and relationships but suffers from scalability issues, as mapping each 3D point to its corresponding object becomes increasingly difficult. In our experiments, it failed to generate reasonable results for prompts with more than four objects; to evaluate larger scenes, we combined multiple objects. LucidDreamer, MVDream, HiFA, and Magic3D perform well on simple prompts but struggle as object count and complexity increase. Their quality significantly declines in highly compositional scenes. In addition to qualitative analysis, we quantitatively compare SOTA methods using CLIP score [33], Pick-A-Pic [19], and a user study on 30 compositional prompts (2–11 objects) in Table 1. In the user study, 20 participants selected the best matches for each prompt based on semantic alignment and quality, with 54% favoring DecompDreamer. DecompDreamer achieves the highest scores across all three metrics, demonstrating its effectiveness in compositional 3D generation.
Disentanglement in Multi-Object 3D Generation Beyond generating compositional 3D scenes, producing high-quality, disentangled individual objects is equally crucial. Figure 5 compares individual objects from DecompDreamer, GALA3D, and GraphDreamer on complex prompts. Both GALA3D and GraphDreamer exhibit poor disentanglement, with GraphDreamer being more prone to object blending due to its stronger focus on relationship modeling. DecompDreamer achieves superior disentanglement through targeted relationship optimization and view-aware object refinement. Additional results are provided in Section A.4 of the appendix.
Table 1: Comparison of CLIP Score, Pick-A-Pic, and User Study results across different text-to-3D methods. All values are reported as percentages.
Method CLIP Score $\uparrow$ Pick-A-Pic $\uparrow$ User Study $\uparrow$
Magic3D 30.77 9.1 0
HiFA 29.55 16.9 2.4
GraphDreamer 28.07 11.6 18.3
GALA3D 31.07 8.6 10.6
LucidDreamer 32.34 15.7 3.3
MVDream 31.71 15.5 11.4
DecompDreamer (Ours) 34.5 22.5 54

### v2

@@ 5 Method
Our objective is to generate high-quality 3D assets for complex text prompts that describe intricate inter-object relationships. Figure 3 shows an overview. Section 5.1 details our structured decomposition using a VLM. Section 5.2 describes the full Composition-Aware optimization pipeline for 3D asset generation.
@@ 5.1 Structured Decomposition with VLMs
Before designing a complex 3D asset, artists typically plan the layout, size, and orientation of objects. Inspired by this workflow, we use a Vision-Language Model (VLM), specifically ChatGPT-4o OpenAI (2024), to construct a compositional scene graph guiding 3D generation. Given a text prompt $t^{g}$, the VLM produces an image $I$ and a scene graph $G=(V,E)$, where $V=\{v_{i}\}_{i=1}^{n}$ represents objects $\{O_{i}\}_{i=1}^{n}$ and $E$ encodes inter-object relationships. The image $I$ provides a visual blueprint to initialize object-level spatial attributes. Each object $O_{i}$ is associated with attributes $\{a_{ij}\}_{j=1}^{m}$ to construct an object prompt $t^{o}_{i}$, Similarly, relationships between object pairs $(O_{i},O_{j})$ are converted into edge prompts $t^{e}_{ij}$, composed from the object names and their described interaction. Using $I$ as reference, we estimate object center coordinates, size, orientation, and azimuth angles, which are used in the View-Aware Disentangled Object Optimization module to supervise individual objects and ensure consistent, disentangled appearance.
Coarse Initialization. Each object $O_{i}$ is represented by a distinct 3D Gaussian cloud, initialized from Point-E Nichol et al. (2022) point clouds and aligned using VLM-inferred spatial properties. For efficient disentangled optimization, we maintain a mapping that tracks Gaussians corresponding to each object, enabling independent refinement of properties such as mean, covariance, and opacity. Further implementation details are in Section A.2.
5.2 Composition-Aware Optimization
@@ 5.2.1 Ensuring Coherent Inter-Object Interactions
Joint Optimization of Related Objects: Given an edge prompt $t^{e}_{ij}$ describing the relationship between objects $O_{i}$ and $O_{j}$, we jointly optimize the Gaussians corresponding to both objects. This ensures that the spatial and semantic relationships between the objects are preserved in the 3D scene. The joint optimization loss is the SDS loss applied to a rendered view containing both objects:
$\mathcal{L}_{joint}(O_{i},O_{j},y^{e}_{ij})=\nabla_{\mathcal{G}_{o_{i}},\mathcal{G}_{o_{j}}}\mathcal{L}_{\text{SDS}}(z;y^{e}_{ij}),$ (2)
where $z$ is the rendered image containing both objects, and $y^{e}_{ij}$ is the text embedding of the edge prompt $t^{e}_{ij}$.
Targeted Optimization for a Single Object: While joint optimization ensures cohesive inter-object relationships, it can sometimes result in undesirable blending, where parts of one object flow into another. To mitigate this, we employ targeted optimization based on the edge prompt, focusing the optimization solely on one object while keeping the other fixed. The goal is to adjust the selected object such that it adheres to the overall relationship without affecting the structural integrity of the other object. The SDS loss for targeted optimization is formulated as:
$\mathcal{L}_{target}(O_{i},y^{e}_{ij})=\nabla_{\mathcal{G}_{o_{i}}}\mathcal{L}_{\text{SDS}}(z;y^{e}_{ij}),$ (3)
@@ 5.2.2 View-Aware Disentangled Object Optimization
DecompDreamer promotes object disentanglement by assigning each Gaussian to a single object, ensuring clear boundaries. However, optimizing only through inter-object relationships can lead to blending and suboptimal refinement of individual components. This section addresses how we optimize object attributes while preserving disentanglement using view-specific and negative prompts.
Optimizing Objects in Compositional Settings. In single-object generation, view-dependent text embeddings Poole et al. (2023) have proven effective for aligning 3D renders with orientation-specific details. However, in compositional settings, a single global camera view often fails to represent all components accurately—leading to conflicting gradients and misalignment between prompts and rendered views. To address this, we propose object-view-dependent supervision. For each object $O_{i}$, we estimate its angular offset $\phi_{i}$ from a canonical front-facing orientation ($0^{\circ}$ azimuth) using the VLM (workflow detailed in Appendix A.8). The corrected azimuth $\psi_{i}=\psi-\phi_{i}$ is used to compute the view-aligned embedding $y^{o}_{\psi_{i}}$. We incorporate this adjustment into our SDS loss, yielding the object-level optimization:
$\mathcal{L}_{\text{obj}}(O_{i},\ y^{o}_{\psi_{i}})=\nabla_{\mathcal{G}_{o_{i}}}\mathcal{L}_{\text{SDS}}(z;\ y^{o}_{\psi_{i}})$ (4)
This refinement ensures that each object is supervised using a view-aligned prompt, improving consistency, orientation accuracy, and geometric coherence in the compositional output.
Negative Prompt for Disentangled Object. To prevent feature blending between related components, we introduce negative prompts during object-specific optimization. For a given object $O_{i}$, the negative prompt is constructed by aggregating the labels of all directly connected objects in the scene graph. This encourages the model to explicitly suppress features from neighboring components, thereby reinforcing object-level separation and preserving clean boundaries.
Figure 4: Qualitative comparison between the proposed DecompDreamer and state-of-the-art text-to-3D generators. More results can be found in Section A.9 of the appendix.
@@ 5.2.3 Overall Optimization
Spatial Error Correction. Before optimizing the full compositional object, we correct spatial misalignments arising from VLM-predicted object centers. To this end, we introduce object-specific translation parameters $T^{o}_{i}=(x_{i},y_{i},z_{i})$ for each object $O_{i}$. These parameters are applied to shift the corresponding Gaussians as $\mathcal{G}_{o_{i}}\leftarrow\mathcal{G}_{o_{i}}+T^{o}_{i}$. To isolate translation refinement, we optimize $T^{o}_{i}$ by minimizing joint edge-level losses:
$\mathcal{L}_{\text{spatial-error-correction}}=\sum_{(i,j)\in E}\mathcal{L}_{\text{joint}}(O_{i},O_{j},y^{e}_{ij})$ (5)
while keeping all other parameters of $O_{i}$ and $O_{j}$ frozen. This adjustment ensures that relationship optimization operates on correctly aligned components, improving the geometric fidelity of the overall composition. After error correction, we begin by modeling inter-object relationships before gradually shifting focus toward individual objects. The optimization is divided into two stages:
Stage One: We emphasize joint modeling of inter-object relationships. For each edge in the scene graph, we apply joint optimization (2), while progressively increasing the weight of individual object optimization (4).
Stage Two: The focus shifts toward refining individual objects while maintaining inter-object consistency. Here, we perform targeted optimization for each object (3) along with individual object optimization (4). This transition allows fine-grained details to be added while preserving relationships established earlier. The effective optimization is:
$\mathcal{L}_{\text{loss}}=\begin{cases}\mathcal{L}_{\text{joint}}(O_{i},O_{j},y^{e}_{ij})+\lambda\left(\frac{t}{T}\right)^{2}\sum\limits_{O\in\{O_{i},O_{j}\}}\mathcal{L}_{\text{obj}}(O,y^{o}_{a_{o}}),&\text{if }t<\gamma T\\[6.0pt] \sum\limits_{O\in\{O_{i},O_{j}\}}\left(\mathcal{L}_{\text{target}}(O,y^{e}_{ij})+\mathcal{L}_{\text{obj}}(O,y^{o}_{a_{o}})\right),&\text{if }t\geq\gamma T\end{cases}$ (6)
where $(i,j)\in E$, $t$ represents the current iteration, $T$ is the total number of iterations, and $\gamma$ controls the transition point between joint and targeted optimization. The full pipeline, including algorithmic details, is provided in Section A.3 of the Appendix.
@@ 6 Experiments
Figure 5: Empirical validation of optimization heuristics. The plots show the average object, scene, and edge losses for a complex 4-object text prompt ("A knight in shining armor…"). (a) GALA3D converges suboptimally, (b) GraphDreamer diverges, while (c) our staged approach (DecompDreamer) converges stably.
Figure 6: Qualitative comparison of object disentanglement in GALA3D, GraphDreamer, and DecompDreamer. Each method displays the full 3D object (left) and its components (right).
Figure 7: Results of an ablation study on DecompDreamer with the text prompt: “A knight in full shining armor galloping on a brown horse, wearing a blue hat, and holding a sword". Additional ablation studies can be found in Figure 8 of the appendix.
We implement our framework in PyTorch, using a rectified flow model (stabilityai/stable-diffusion-3.5-medium) for guidance and representing scenes with 3D Gaussian Splatting. Key hyperparameters, including guidance scale, Gaussian count, and optimization iterations, are detailed in Appendix A.1. Our experiments validate our approach through an analysis of loss dynamics (Section 6.1), comparisons with state-of-the-art methods (Section 6.2), and ablation studies (Section 6.3). A typical two-object scene is generated in 90 minutes on a single A100 GPU; a full runtime comparison with baselines is provided in Appendix A.4.
6.1 Quantifying Gradient Conflicts
To empirically validate our taxonomy, we analyze the loss dynamics of representative methods in Figure 5. The results visually confirm the predicted failure modes of prior heuristics and demonstrate the stability of our staged approach. The loss curves for GALA3D (Figure 5a) illustrate the predicted stable but suboptimal convergence. After an initial descent, both object and scene losses hit a floor and begin to rise again, eventually settling at a high error value. This trajectory reveals the inherent conflict between optimizing object fidelity and adhering to rigid layout priors, leading the system to converge to a poor local minimum. The trajectory for GraphDreamer (Figure 5b) provides a clear illustration of optimization divergence. The object and edge losses rapidly increase, indicating the optimizer is immediately overwhelmed by a combinatorial explosion of conflicting gradients. This creates a chaotic "tug-of-war," visually evident as the loss curves cross while trending upwards, which suggests that an improvement in one objective comes at the direct expense of the other, leading to a failed optimization. In stark contrast, the loss curves for DecompDreamer (Figure 5c) demonstrate the effectiveness of our staged curriculum. The optimization proceeds in two clean phases. In Stage 1, the relational loss converges sharply while the object loss decreases more slowly. In Stage 2, the relational loss remains stable, proving the structure is preserved, while the object loss begins its own rapid descent. This plot is definitive proof of temporal decoupling, which avoids the gradient conflicts that plague other heuristics and ensures a stable convergence.
@@ 6.2 Comparison to State-of-the-art Methods
We evaluate DecompDreamer against state-of-the-art text-to-3D methods, including Gala3D, GraphDreamer, MVDream, LucidDreamer, Magic3D, and Trellis, on a series of complex compositional prompts. Although CompGS reports competitive results, we were unable to include it due to the lack of publicly available code at the time of our experiments. Qualitatively, as shown in Figures 1 and 4, DecompDreamer excels at modeling intricate inter-object relationships and scales effectively to a large number of objects. In contrast, methods like GALA3D can generate large environments but struggle to model specific interactions, while GraphDreamer effectively models relationships but fails to scale beyond a few objects.
Our quantitative evaluation, summarized in Table 1, confirms these findings through automated metrics and a user study (20 participants, 30 prompts, 2-11 objects; see Appendix A.5). A clear overall user preference for DecompDreamer (61%) becomes more pronounced as scene complexity increases. For prompts with more than three objects, user preference surges to 73.0%, and our Text-to-3D Alignment score (86.6%) nearly doubles that of the nearest competitor. While baselines can achieve competitive scores in simple scenes ($\leq$ 3 objects), their performance collapses as relational complexity increases—for instance, GraphDreamer’s alignment score plummets from 50.0% to 16.5%. Critically, even in these simpler cases, users still prefer DecompDreamer (53.0%), demonstrating its decisive advantage regardless of object count. This analysis confirms that our framework is uniquely capable of handling both scale and relational complexity, a key failure point for prior work.
Disentanglement in Multi-Object 3D Generation Beyond compositional fidelity, producing well-separated object components is critical. Figure 6 shows that Gala3D and GraphDreamer exhibit poor disentanglement, with GraphDreamer prone to blending due to its stronger focus on relationship modeling. In contrast, DecompDreamer maintains clear boundaries through targeted relationship optimization and view-aware object refinement. Additional results are provided in Appendix A.7.
Table 1: Comparison of CLIP Score, Pick-A-Pic, Text-to-3D Alignment, and User Study across different object count ranges. All values are percentages.
Method CLIP Score $\uparrow$ Pick-A-Pic $\uparrow$ Text-to-3D Alignment $\uparrow$ User Study $\uparrow$ # of Objects $\leq$3 >>3 All $\leq$3 >>3 All $\leq$3 >>3 All $\leq$3 >>3 All Magic3D 32.9 29.5 31.8 12.0 8.0 11.0 59.6 19.6 46.7 0.0 0.0 0.0 GraphDreamer 32.6 19.4 28.4 17.0 13.0 16.0 50.0 16.5 39.2 22.0 14.0 19.0 GALA3D 30.3 30.1 30.2 11.0 6.0 9.0 69.3 51.8 63.7 12.0 6.0 10.0 LucidDreamer 32.8 33.5 33.0 15.0 21.0 17.0 59.6 21.3 47.3 6.0 0.0 3.0 MVDream 32.0 32.3 32.1 16.0 15.0 16.0 48.9 16.1 38.4 0.0 0.0 0.0 Trellis 28.7 27.7 28.4 8.0 5.0 7.0 41.0 19.7 34.1 7.0 7.0 7.0 Ours 34.7 34.6 34.5 21.0 32.0 24.0 67.2 86.6 73.4 53.0 73.0 61.0